DLP ACTIVITY #4

Direction: Calculate the following word problems involving limiting and excess reagents, theoretical

and percent yield. Show your solution.

1. Silver metal reacts with sulfur to form silver sulfide according to the following reaction:

2Ag+S→Ag2S

a. Identify the limiting reagent if 50.0 g Ag reacts with 10.0 g S.

b. What is the theoretical yield in g of Ag2S produced from the reaction?

c. What is the amount in g of the excess reactant expected to remain after the reaction?

d. When the reaction occurred, the amount of Ag2S obtained was 45.0 g. What is the percent yield

of the reaction?

2. Calculate the mass of magnesium oxide possible if 2.40 g Mg reacts with 10.0 g O2.

Mg+O2 →MgO

a. What is the limiting reagent?

b. What is the theoretical yield?

c. What is the excess reagent and by how much?

3. Given the following reaction: Zn + CuCl2 → ZnCl2 + Cu

a. How many grams of ZnCl2 will be produced from 7.00 g of Zn and 8.00 g of CuCl2?

b. What is the limiting reagent?



1
Expert's answer
2021-12-13T22:37:01-0500

2Ag(s) + S(s) → Ag2S(s)


For one mole of sulphur there should be two moles of silver for complete reaction.

Moles of silver = 50/108 = 0.46 moles

Moles of elemental sulphur = 10/32 = 0.3125 moles

For 0.3125 moles of sulphur there must be 0.6250mole of silver.

So silver is limiting Reagent.